# references.bib

### v1

@@ 1 Introduction
when self-loops are excluded. Here $D$ is the $n\times n$ diagonal matrix containing node hyperdegrees (the number of hyperedges in which the node occurs). Throughout this work, we will consider co-occurrence projections both with and without hyperdegrees on the diagonal.
@@
Definition 1.3 (Gram Mates [kim_gram_2023]).
@@ 2 Enumerating sets of twins
4:     Choose edge $e=(u,v)\in V_{e}$ s.t. $W_{uv}>0$
5: // Initialize stack with combinations
@@
24:              for $\delta\in{\eta^{\prime}_{e}\choose W^{\prime}_{e}}$ do
@@
The algorithm proceeds as follows. First, we construct the factor graph $F=(V_{e},V_{c},\eta_{e})$ from $W$. We then choose an edge-node $e$ with positive $W_{uv}$.  [^2] In practice, we choose this edge in the fastest way possible, thus making the order of execution arbitrary. One can impose an ordering on the satisfaction of edges to achieve a specific goal. For example, to minimize the size of the search tree, we can satisfy fully-determined edges first by always choosing the edge that minimizes $c(e)={|\eta_{e}|\choose W_{uv}}$. When $c(e)=1$, all of the clique-nodes attached to edge-node $e$ in the factor graph must be included in a hypergraph that satisfies the constraints of $W$. By satisfying edges with $c(e)=1$ immediately, we avoid adding a layer of depth later in the search tree when it may be wider, thus adding unnecessary nodes, and start from the smallest possible problem. However, in the worst case computing this constraint requires a loop over the pairwise edges at every node in the search tree, which may not always be practical. Every combination of $W_{uv}$ clique-node-neighbors from $\eta_{e}$, each corresponding to a partial hypergraph that satisfies $e$, is added to a stack (or queue) before entering the main loop.
@@ 2.2 Computational Complexity
Once the factor graph has been constructed, we then choose an edge $e$ in constant time and add ($\eta_{e}\choose W_{uv}$ nodes to the stack or queue, corresponding to the possible combinations of cliques connected to $e$ that are compatible with $W$. This is the main work that needs to be done at every node of the search tree.
We now turn to the size of the search tree. A trivial upper bound on the depth of the tree is $\mathcal{O}{(|E|)}$, meaning there is one layer for every edge in the projection. In practice, especially for sparse projections, we would expect the depth to be smaller, since some edges will be satisfied incidentally through the satisfaction of earlier edges.
The maximum number of items that can be added to the stack is an upper bound on the width of each layer. This number is bounded by the maximum number of possible sets of cliques that could satisfy any edge $e$, computed on the full factor graph:
$C=\max\limits_{e}{{|\eta_{e}|\choose W_{uv}}}.$
This gives a trivial worst-case time of $\mathcal{O}{(C^{|E|})}$, the case where every layer of the tree has the same width. A tighter bound is simply the product of the widths:
$\mathcal{O}(\prod\limits_{e}{|\eta_{e}|\choose W_{uv}}).$
Later we will show numerical results regarding the relationship between this worst-case running time and properties of the input, as well as wall-clock running time in our C++ implementation [larock_TwinSearch_2023].
@@ 3.1 Combinatorial Upper Bounds
((${{n\choose k}\choose\mathcal{M}_{k}}.$
@@
(${\mathcal{C}_{k}\choose\mathcal{M}_{k}}.$
We know that $\mathcal{C}_{k}\leq{n\choose k}$, therefore this second bound is always better except in the case where every possible clique on $k$ nodes exists in the projection. [^3] We note that in this analysis the cliques are binary while the projection is weighted. Thus having all possible cliques in the projection does not necessarily fix the set of twins.
@@
Figure 2: Example describing how binomial coefficients overcount the number of hypergraphs associated to a projection in the 3-uniform case. Given the input projection on the left, when $\mathcal{M}_{k}=4$ hyperedges of size $k=3$, there is 1 extra hypergraph counted by the binomial coefficient ($\mathcal{C}_{k}\choose\mathcal{M}_{k}$ that does not have the input projection as its node co-occurrence projection matrix.
@@
$\prod_{k\textrm{ s.t. }\mathcal{M}_{k}>0}{\mathcal{C}_{k}\choose\mathcal{M}_{k}}.$
Each term in this product is susceptible not only to the overcounting described above, but also to overcounting by not respecting the weights in $W$ in other ways. For example, if we assume there is an edge in $W$ with weight 1 that participates in at least 1 triangle, the above product counts hypergraphs where that edge would appear both as a 2-hyperedge and in a 3-hyperedge, which would result in a weight larger than 1.
The question remains how many of those possible hypergraphs actually correspond to the input projection $W$? More complex counting arguments may be fruitful, however we elect to approach the problem numerically, using the TwinSearch algorithm to investigate the relationship between the weighted edge structure of a node co-occurrence projection and its corresponding set of twins.
@@ 3.3 Case Study: Exhaustive Search
We now turn to numerically analyzing sets of twins. We will focus our analysis on a modified Erdős-Rényi model for hypergraphs in which we choose $m$ hyperedges uniformly without replacement from ($n\choose k$ possible hyperedges, rejecting hypergraphs where some node $u$ is not included in any of the $m$ sampled hyperedges (i.e., $u$ is a singleton), guaranteeing that the minimum hyperdegree of any node is 1. We choose to impose this constraint in this context because for our algorithm there is no difference between a hypergraph on $n$ nodes that includes $n_{s}$ singletons and a hypergraph on $n-n_{s}$ nodes, since all entries in the projection matrix involving any of the $n_{s}$ singleton nodes will be 0.
In this section we will focus in detail on projections of hypergraphs from the $3$-uniform ensemble on $6$ nodes, denoted $G_{3}(6,m)$. We focus on these parameters because it is practical to enumerate all simple hypergraphs and their node co-occurrence projections across all values of $m$ between 2 and (${6\choose 3}=20$.
@@
In Figure 3 we study only $3$-uniform hypergraphs found by the algorithm. The middle plot of Figure 3 shows the number of $3$-uniform twins as a percentage of the number of unique projections for each $m$ (shown in the left-hand plot). If every projection was realized by exactly one unique hypergraph, both lines would be flat at 100%, as we see outside of the range from $m=4\dots 17$ for all twins (green) and outside the range $m=7\dots 13$ for isomorphism classes (purple). The total number of twins reaches a peak of 15% larger than the number of projections at $m=10$, while the total isomorphism classes is maximum for $m=9,10,11$ at around 2.5% of the number of projections.
Figure 4: Relationships between number of hyperedges $m$ and the distribution of various quantities of interest, including: (a) the number of non-uniform twins; (b) the number of isomorphism classes among the set of twins; (c) the number of unique hyperedge size distributions; (d) the number of cliques in the projection; (e) worst case search tree size; (f) elapsed running time; and pairs of exact (g) and isomorphic (h) Gram Mates. Each boxplot corresponds to the distribution of the quantity per projection from hypergraphs with $m$ hyperedges. Boxes span from 1st to 3rd quartiles, orange lines show median values, and whiskers represent minimum and maximum values. The blue lines in (g-h) show the number of projections with at least one pair of Gram Mates.
@@ 3.4 Non-uniform Twins
We now expand our analysis to sets of twins including non-uniform hypergraphs, where the hyperedge sizes are constrained only by the clique size distribution in the projection (equiv. the clique-nodes in the factor graph). Figure 4 shows relationships between the number of hyperedges $m$ and distributions of various statistics among the full sets of twins associated to node co-occurrence projections. The distributions are shown as box and whisker plots, where the limits of the box show the 1st and 3rd quartiles, the orange lines show the median, and the whiskers extend to the minimum and maximum values in the data.
The top row shows the distributions of the number of non-uniform twins (Figure 4(a)) and isomorphism classes (b). The number of twins grows as an exponent of $m$ and the distributions are not noticeably different, meaning that the majority of twins are unique up to isomorphism.
@@
Returning to Figure 4, plot (d) shows the distribution of the total number of cliques of sizes $2,\dots 6$ in the input projections. With $m=6$, there are already some projections that have the maximum number of cliques, which we can compute as
$\sum_{k\in 3,\dots,6}{6\choose k}=42\textrm{,}$
and after $m=16$ all possible cliques are included in every projection, highlighting the importance of the weighted structure of the projection.
Figure 4(e) shows the distribution of the worst case search tree size (see Section 2.2 for details). As the number of hyperedges grows, the worst case complexity increases exponentially. Figure 4(f) shows the elapsed runtime in our C++ implementation [larock_TwinSearch_2023]. The runtime grows more slowly than the worst case search size, but still at an exponential rate with respect to $m$.
@@
Taken together, the results in Figures 4 and 5 exemplify the diversity within the set of non-uniform twins corresponding to the node co-occurrence projection in terms of the number of hyperedges, the distribution of hyperedge sizes, and the node hyperdegrees.
@@ 3.6 Sampled k-uniform Search
This sampling procedure has two important implications. First, as stated above, it trivially guarantees that every sampled projection corresponds to at least one $k$-uniform hypergraph. Second, it implies that our results are likely to be over-counts, since we may sample the same projection multiple times either by sampling the same hypergraph repeatedly or by sampling different hypergraphs that have the same projection. This is especially likely when the binomial coefficient governing the number of possible hypergraphs is small relative to the number of samples. Similarly, our counts are also potentially inflated by isomorphism between projections, since two projections that are isomorphic will have the same number of twins. While overcounting is not ideal, the trade-off is that we can easily do the sampling and computations in parallel over a wider range of parameters than if we need to pairwise compare every sampled projection to all previous projections to ensure uniqueness.
@@
In the right column of Figure 7, we expand the columns of the heatmaps for $n\in 7,8,9$ for $k=3$ and $n\in 7,8$ for $k=4$. The plots show the average and standard deviation of the number of isomorphism classes per sampled projection across the entire hyperedge space for simple hypergraphs for each $k,n$ pair, i.e., when ((${{n\choose k}\choose m}>0$.
In general, as the number of nodes increases, so does the maximum number of twins per hypergraph. While the shape of the curves is similar, the order of magnitude of the average number of isomorphism classes increases substantially with $k$ and $n$, reaching averages above $10^{3}$ for $k=3,n=9$ and $k=4,n=8$. Further, the symmetry of the curves over the full range of $m$ indicates that increasing the density of the hypergraph increases the average size of the $k$-uniform set of non-isomorphic twins so long as it increases the number of possible hypergraphs, that is, while the binomial coefficient is increasing. This suggests that, similar to binomial coefficients in general, for a number of hyperedges $m$ the set of complements with (${n\choose k}-m$ hyperedges is likely to have similar properties. We leave further analysis of this dual relationship for future work.

### v2

@@ 1 Introduction
when self-loops are excluded. Here $D$ is the $n\times n$ diagonal matrix containing node hyperdegrees (the number of hyperedges in which the node occurs). Throughout this work, we will consider co-occurrence projections both with and without hyperdegrees on the diagonal. [^2] We note that throughout this paper we will use $W_{uv}$, $W_{e}$ (with index $e=(u,v)$ assumed), and simply $W$ interchangeably depending on context. Each refers to a node co-occurrence projection matrix.
@@
Definition 1.3 (Gram Mates ).
@@ 2 Enumerating sets of twins
4:     Choose edge $e\in V_{e}$ s.t. $W_{e}>0$
5: // Init. stack with combinations
@@
24:              for $\delta\in\binom{\eta^{\prime}_{e}}{W^{\prime}_{e}}$ do
@@
The algorithm proceeds as follows. First, we construct the factor graph $F=(V_{e},V_{c},\eta_{e})$ from $W$. We then choose an edge-node $e$ with positive $W_{e}$.  [^3] In practice, we choose this edge in the fastest way possible, thus making the order of execution arbitrary. One can impose an ordering on the satisfaction of edges to achieve a specific goal. For example, to minimize the size of the search tree, we can satisfy fully-determined edges first by always choosing the edge that minimizes $c(e)=\binom{|\eta_{e}|}{W_{uv}}$. When $c(e)=1$, all of the clique-nodes attached to edge-node $e$ in the factor graph must be included in a hypergraph that satisfies the constraints of $W$. By satisfying edges with $c(e)=1$ immediately, we avoid adding a layer of depth later in the search tree when it may be wider, thus adding unnecessary nodes, and start from the smallest possible problem. However, in the worst case computing this constraint requires a loop over the pairwise edges at every node in the search tree, which may not always be practical. Every combination of $W_{e}$ clique-node-neighbors from $\eta_{e}$, each corresponding to a partial hypergraph that satisfies $e$, is added to a stack (or queue) before entering the main loop.
@@ 2.2 Computational Complexity
Once the factor graph has been constructed, we then choose an edge $e$ in constant time and add $\binom{\eta_{e}}{W_{e}}$ nodes to the stack or queue, corresponding to the possible combinations of cliques connected to $e$ that are compatible with $W$. This is the main work that needs to be done at every node of the search tree.
We now turn to the size of the search tree. A trivial upper bound on the depth of the tree is $\mathcal{O}{(|E|)}$, meaning there is one layer for every edge in the projection. In practice, especially for sparse projections, we expect the depth to be smaller, since some edges will be satisfied incidentally through the satisfaction of earlier edges.
The maximum number of items that can be added to the stack is an upper bound on the width of each layer. This number is bounded by the maximum number of sets of cliques that could satisfy any edge $e$, computed on the full factor graph:
$C=\max\limits_{e}{\binom{|\eta_{e}|}{W_{e}}}.$
This gives a trivial worst-case search tree size of $\mathcal{O}{(C^{|E|})}$, the case where every layer of the tree has the same width. A tighter bound on the worst case search tree size is simply the product of the widths at every edge:
$\mathcal{O}(\prod\limits_{e}{\binom{|\eta_{e}|}{W_{e}}}).$ (1)
Later we will show numerical results regarding the relationship between this worst-case running time and properties of the input, as well as wall-clock running time in our C++ implementation (see Figure 6) [larock_TwinSearch_2023].
@@ 3.1 Combinatorial Upper Bounds
$\binom{\binom{n}{k}}{\mathcal{M}_{k}}.$
@@
$\binom{\mathcal{C}_{k}}{\mathcal{M}_{k}}.$
We know that $\mathcal{C}_{k}\leq\binom{n}{k}$, therefore this second bound is always better except in the case where every possible clique on $k$ nodes exists in the projection. [^4] We note that in this analysis the cliques are binary while the projection is weighted. Thus having all possible cliques in the projection does not necessarily fix the set of twins.
@@
Figure 2: Example describing how binomial coefficients overcount the number of hypergraphs associated to a projection in the 3-uniform case. Given the input projection on the left, when $\mathcal{M}_{k}=4$ hyperedges of size $k=3$, there is 1 extra hypergraph counted by the binomial coefficient $\binom{\mathcal{C}_{k}}{\mathcal{M}_{k}}$ that does not have the input projection as its node co-occurrence projection matrix.
@@
$\prod_{k\textrm{ s.t. }\mathcal{M}_{k}>0}\binom{\mathcal{C}_{k}}{\mathcal{M}_{k}}.$ (2)
We will refer to the product of binomials in Equation 2 as the clique approximation to the size of the set of twins. Each term in this product is susceptible not only to the overcounting described above, but also to overcounting by not respecting the weights in $W$ in other ways. For example, if we assume there is an edge in $W$ with weight 1 that participates in at least 1 triangle, the above product counts hypergraphs where that edge would appear both as a 2-hyperedge and in a 3-hyperedge, which would result in a weight larger than 1.
The question remains how many of those possible hypergraphs actually correspond to the input projection $W$? More complex counting arguments may be fruitful, however we elect to approach the problem numerically, using the TwinSearch algorithm to investigate the relationship between the weighted edge structure of a node co-occurrence projection and its corresponding set of twins. Later, in Figure 6, we will show the relationship between the true number of twins with a given hyperedge size distribution $\mathcal{M}_{k}$ corresponding to a projection with a given clique size distribution $\mathcal{C}_{k}$ and the approximation given in Equation 2.
@@ 3.3 Case Study: Exhaustive Search
We now turn to numerically analyzing sets of twins. We will focus our analysis on a modified Erdős-Rényi model for hypergraphs in which we choose $m$ hyperedges uniformly without replacement from $\binom{n}{k}$ possible hyperedges, rejecting hypergraphs where some node $u$ is not included in any of the $m$ sampled hyperedges (i.e., $u$ is a singleton), guaranteeing that the minimum hyperdegree of any node is 1. We choose to impose this constraint in this context because for our algorithm there is no difference between a hypergraph on $n$ nodes that includes $n_{s}$ singletons and a hypergraph on $n-n_{s}$ nodes, since all entries in the projection matrix involving any of the $n_{s}$ singleton nodes will be 0.
In this section we will focus in detail on projections of hypergraphs from the $3$-uniform ensemble on $6$ nodes, denoted $G_{3}(6,m)$. We focus on these parameters because it is practical to enumerate all simple hypergraphs and their node co-occurrence projections across all values of $m$ between 2 and $\binom{6}{3}=20$.
@@
In Figure 3 we focus on $3$-uniform hypergraphs found by the algorithm. The middle plot shows the number of $3$-uniform twins as a percentage of the number of unique projections for each $m$ (shown in the left-hand plot). If every projection was realized by exactly one unique hypergraph, both lines would be flat at 100%, as we see outside of the range from $m=4\dots 17$ for all twins (green) and outside the range $m=7\dots 13$ for isomorphism classes (purple). The total number of twins reaches a peak of 15% larger than the number of projections at $m=10$, while the total isomorphism classes is maximum for $m=9,10,11$ at around 2.5% of the number of projections.
Figure 4: Relationships between number of hyperedges $m$ and the distribution of various quantities of interest, including: (a) the number of non-uniform twins; (b) the number of isomorphism classes among the set of twins; (c) the number of unique hyperedge size distributions; (d) the number of cliques in the projection; (e) worst case search tree size; (f) elapsed running time (note: vertical axis is cut off in log scale due to 0s corresponding to sub-millisecond runtimes); and pairs of exact (g) and isomorphic (h) Gram Mates. Each boxplot corresponds to the distribution of the quantity per projection from hypergraphs with $m$ hyperedges. Boxes span from 1st to 3rd quartiles, orange lines show median values, and whiskers represent minimum and maximum values. The blue lines in (g-h) show the number of projections with at least one pair of Gram Mates.
@@ 3.4 Non-uniform Twins
We now expand our analysis to sets of twins including non-uniform hypergraphs, where the hyperedge sizes are constrained only by the clique size distribution in the projection (equiv. the clique-nodes in the factor graph), i.e., hyperedges can in principle range in size from $2,\dots,n$. Figure 4 shows relationships between the number of hyperedges $m$ and distributions of various statistics among the full sets of twins associated to unique up to isomorphism node co-occurrence projections computed from hypergraphs in $G_{3}(6,m)$. The distributions are shown as box and whisker plots, where the limits of the box show the 1st and 3rd quartiles, the orange lines show the median, and the whiskers extend to the minimum and maximum values in the data.
The top row shows the distributions of the number of non-uniform twins (Figure 4(a)) and isomorphism classes (b). The number of twins grows monotonically with $m$ and the distributions are not noticeably different, implying that a substantial proportion of twins are unique up to isomorphism.
@@
Figure 6: Left: Number of twins against worst case search tree size (see Eq. 1) for unique up to isomorphism projection matrices from $G_{3}(6,m)$ (color in the first 3 plots corresponds to $m$). Middle: Numerical evidence of overcounting using the clique approximation (see Eq. 2, Fig. 2, and surrounding discussion). The second plot shows results for every observed combination of clique size distribution $\mathcal{C}_{k}$ and hyperedge size distribution $\mathcal{M}_{k}$, while the third plot shows the average over all combinations for each projection. Right: Kendall’s $\tau$ of rankings of projections by worst case search tree size (Eq. 1, orange triangles) and the sum of the clique approximations for every combination of clique and hyperedge size distribution (Eq. 2, blue circles), both compared against the ranking of projections by their total number of non-uniform twins.
@@
Returning to Figure 4, plot (d) shows the distribution of the total cliques of sizes $3,\dots,6$ in the input projections. With $m=6$, there are already some projections that have the maximum number of cliques, which we can compute as:
$\sum_{k\in 3,\dots,6}\binom{6}{k}=42\textrm{,}$
and after $m=16$ all possible cliques are included in every projection, highlighting the importance of the weighted structure of the matrices.
Figure 4(e) shows the distribution of the worst case search tree size (Equation 1). As the number of hyperedges grows, the worst case complexity increases exponentially. Figure 4(f) shows the elapsed runtime in our C++ implementation [larock_TwinSearch_2023]. The runtime grows more slowly than the worst case search size, but still at an exponential rate with respect to $m$.
@@
Taken together, the results in Figures 4 and 5 exemplify the diversity within the set of non-uniform twins corresponding to the node co-occurrence projection in terms of the number of hyperedges, the distribution of hyperedge sizes, node hyperdegrees, and Gram Mates. Our analysis so far shows that many key characteristics correlate with the density of hyperedges $m$. Notably, the worst case search tree size appears to correlate strongly with both the value of $m$ and, comparing across Figure 4(a) and (e), the total number of non-uniform twins. This observation leads to our final analysis in this section, where we ask whether we can use approximations that do not require enumerating sets of twins to predict the number of twins associated to an individual projection.
In the leftmost plot of Figure 6, we show the worst case search tree size (Equation 1) against the total number of twins for each projection. We see a clear correlation between the two variables, meaning that a larger set of twins implies a larger search tree. This is on the one hand intuitive, since each twin is a leaf in the search tree, thus the number of twins is a trivial lower bound on the tree size. However, as can be seen from the order of magnitude of the worst case search tree size, there are large numbers of search tree nodes, both non-twin leaves and intermediate nodes, whose quantity could in principle have been less directly correlated to the true twin set size. What is powerful about the relationship between the number of twins and the search tree size is that it allows for the possibility of comparing projection matrices using just Equation 1, which can be computed based only on the node co-occurrence matrix and its factor graph, without enumerating the full set of twins with TwinSearch.
We have also described another approximation for the number of twins, the clique approximation introduced in Section 3.1. In the middle two plots of Figure 6, we show the relationship between the number of twins and Equation 2. The second plot shows results for all combinations of clique size distribution $\mathcal{C}_{k}$ and hyperedge size distribution $\mathcal{M}_{k}$ found in the data, while the third plot shows the average computed over all observed combinations of the two distributions per projection. Broadly, the approximation overcounts by orders of magnitude, but similarly to the worst case search tree size, there is an apparent correlation between the true number of twins and the approximation.
We now have evidence that our two approximations correlate with the true number of twins. However, the extent to which this information is practically useful remains to be seen. As a step in this direction, we use a simple rank correlation approach to understand whether these two approximations might be used to compare node co-occurrence projection matrices in terms of their number of twins, without running TwinSearch.
In the rightmost plot of Figure 6, we show Kendall’s $\tau$ rank correlations based on rankings of projection matrices. For each input hyperedge size $m$ such that there are at least 2 unique up to isomorphism node co-occurrence projections, we compute three rankings. First, we rank the projections by their total number of non-uniform twins; second, we rank by the the worst case search tree size (Eq. 1); third, by the clique approximation (Eq. 2), summed over all observed combinations of $\mathcal{C}_{k}$ and $\mathcal{M}_{k}$. We then compute Kendall’s $\tau$ rank correlation coefficients between the number of twins ranking and the two other rankings based on the approximations.
The rank correlation statistics for both approximations are generally positive and significant (below $0.05$), with some exceptions at the lowest and highest values of $m$, where there are relatively few projections (numerical coefficients and p-values are available in Appendix A, Table 1). While neither ranking is fully concordant with the number of twins, which would be the ideal outcome, the approximations do appear to provide some meaningful information about the relative number of twins among different projection matrices. This is promising evidence that it may be feasible to compare node co-occurrence projection matrices with respect to their sets of twins without enumerating them, however we leave further exploration of this possibility for future work.
@@ 3.6 Sampled k-uniform Search
This sampling procedure has two important implications. First, as stated above, it trivially guarantees that every sampled projection corresponds to at least one $k$-uniform hypergraph. Second, it implies that our results are likely to be overcounts, since we may sample the same projection multiple times either by sampling the same hypergraph repeatedly or by sampling different hypergraphs that have the same projection. This is especially likely when the binomial coefficient governing the number of possible hypergraphs is small relative to the number of samples. Similarly, our counts are also potentially inflated by isomorphism between projections, since two projections that are isomorphic will have the same number of twins. While overcounting is not ideal, the trade-off is that we can easily do the sampling and computations in parallel over a wider range of parameters than if we need to pairwise compare every sampled projection to all previous projections to ensure uniqueness.
@@
In the right column of Figure 8, we expand the columns of the heatmaps for $n\in 7,8,9$ for $k=3$ and $n\in 7,8$ for $k=4$. The plots show the average and standard deviation of the number of isomorphism classes per sampled projection across the entire hyperedge space for simple hypergraphs for each $k,n$ pair, i.e., when $\binom{\binom{n}{k}}{m}>0$.
In general, as the number of nodes increases, so does the maximum number of twins per hypergraph. While the shape of the curves is similar, the order of magnitude of the average number of isomorphism classes increases substantially with $k$ and $n$, reaching averages above $10^{3}$ for $k=3,n=9$ and $k=4,n=8$. Further, the symmetry of the curves over the full range of $m$ indicates that increasing the density of the hypergraph increases the average size of the $k$-uniform set of non-isomorphic twins so long as it increases the number of possible hypergraphs, that is, while the binomial coefficient is increasing. This suggests that, similar to binomial coefficients in general, for a number of hyperedges $m$ the set of complements with $\binom{n}{k}-m$ hyperedges is likely to have similar properties. We leave further analysis of this dual relationship for future work.